# Multimodal Distributions for Circular Axial Data

### v1

@@ Abstract
The family of circular distributions based on non-negative trigonometric sums (NNTS), developed by Fernández-Durán (2004), is highly flexible for modeling datasets exhibiting multimodality and/or skewness. In this article, we extend the NNTS family to axial data by identifying conditions under which the original NNTS family is suitable for modeling undirected vectors. Since the estimation is performed using maximum likelihood, likelihood ratio tests are developed for characteristics of the density function such as uniformity and symmetry. The proposed methodology is applied to real datasets involving orientations of rocks, animals, and plants.
@@ 1 Introduction
Axial data, defined as data representing undirected vectors, are relevant in many applications across different disciplines, notably in Paleomagnetism for the statistical analysis of Earth’s magnetic field directions recorded during rock formation. Axial data also appear in analyses of astronomical body orientations, protein structure axes, and certain anatomical structures in animals and humans, such as the posterior corneal curvature of the eye. In Ecology, axial data assist in tracking animal magnetic orientation, and in Environmental Science, they aid the study of vegetation orientation patterns, their relation to climate change, and associated effects on wind direction.
Arnold and Sengupta (2011) describe various methods to generate densities for an axial random variable, $\theta\in[0,\pi)$, from either a linear random variable $X$ or a circular random variable $\phi\in[0,2\pi)$. The first method involves wrapping a linear random variable: $\theta=X\mod\pi$. The second, called adding, defines the axial density as the sum of a circular density at $\phi$ and $\phi+\pi$, equivalent to wrapping a circular variable: $\theta=\phi\mod\pi$. The third method, doubling, defines $\theta=2\phi$, which Arnold and Sengupta (2011) argue is generally inadequate. The fourth method uses polar transformation of a bivariate linear vector $(X,Y)$, defining $\theta=\tan^{-1}(Y/X)+\pi/2$.
Common models for axial data include the axial von Mises (AvM), axial wrapped Cauchy (AWC), and sine-skewed distributions. The AvM model (Arnold and SenGupta, 2006), derived using the adding method, has the density:
@@
$f_{AWC}(\theta)=\frac{1-\rho^{4}}{\pi(1+\rho^{4}-2\rho^{2}\cos(2\theta))},\quad-\frac{\pi}{2}\leq\theta<\frac{\pi}{2},$
where $0\leq\rho\leq 1$. Additionally, Arnold and SenGupta (2011) introduced the angular central Gaussian (AACG) model, which generalizes the AWC by using a polar transformation of a bivariate Gaussian vector:
$f_{AACG}(\theta)\propto\left(1+\alpha\sin\left(2\left(\theta+\frac{\pi}{4}-\beta\right)\right)\right)^{-1},\quad 0\leq\theta<\pi,$
with parameters $\alpha$ and $\beta$. Abe et al. (2012) proposed sine-skewed (SineSk) axial models, introducing asymmetry through a perturbation term:
$f_{SineSk}(\theta)=(1+\lambda\sin(k\theta))f_{0}(\theta),\quad-\frac{\pi}{2}\leq\theta<\frac{\pi}{2},$
where $f_{0}(\theta)$ is the base symmetric axial density, $k=2$, and $-1\leq\lambda\leq 1$. When $\lambda=0$, the SineSk model reduces to the symmetric base model. Multimodal axial densities may also be constructed as mixtures (e.g., mixtures of axial von Mises distributions), though this approach complicates numerical parameter estimation when the number of modes increases.
@@ 2.1 Definition
$f_{NNTS}(\phi\mid\underline{c})=\left|\left|\sum_{k=0}^{M}c_{k}e^{ik\phi}\right|\right|^{2}=\sum_{k=0}^{M}\sum_{m=0}^{M}c_{k}\bar{c}_{m}e^{i(k-m)\phi},$ (1)
@@
$\sum_{k=0}^{M}||c_{k}||^{2}=\frac{1}{2\pi},$
where $||c_{k}||^{2}=c_{re,k}^{2}+c_{im,k}^{2}$. This implies that $c_{im,0}=0$, making $c_{0}$ a positive real number.
@@
$c_{k}^{*}=c_{k}\text{ for even }k\quad\text{and}\quad c_{k}^{*}=-c_{k}\text{ for odd }k.$
@@
$f_{NNTSaxial}(\theta\mid\underline{v})=\left|\left|\sum_{k=0}^{M}v_{k}e^{i2k\theta}\right|\right|^{2}=\sum_{k=0}^{M}\sum_{m=0}^{M}v_{k}\bar{v}_{m}e^{i2(k-m)\theta}.$ (3)
This corresponds to doubling the original circular angle $\phi$, a common technique for analyzing axial data. For this to be a valid density, the parameters must satisfy
$\sum_{k=0}^{M}||v_{k}||^{2}=\frac{1}{\pi},$
implying $v_{0}$ must be a non-negative real number. Alternatively, the NNTS axial density can be expressed as:
@@
with $\sum_{k=0}^{M}||v_{k}||^{2}=1$, placing parameters on the unit complex hypersphere $CS^{M}=\{\underline{v}:\sum_{k=0}^{M}||v_{k}||^{2}=1\}$.
The number of free parameters in $\underline{v}$ equals $2M$, and $M$ determines the maximum number of modes for the NNTS axial density. The support of the axial random variable can be $(0,\pi]$ due to $f_{NNTSaxial}(\theta\mid\underline{v})=f_{NNTSaxial}(\theta\pm\pi\mid\underline{v})$, although the interval $[-\frac{\pi}{2},\frac{\pi}{2})$ is also common in literature. Note that NNTS axial models are nested; models with smaller $M^{*}$ are special cases of those with larger $M^{**}$, facilitating likelihood ratio tests. Specifically, the case $M=0$ corresponds to the uniform axial density.
The trigonometric moments, $E(e^{ir\theta})$, for $r$ a non-negative integer, of the NNTS axial distribution in Equation (4), are given by:
@@
$E(e^{ir\theta})=\sum_{m=r}^{M}v_{m-r}\bar{v}_{m}$ (6)
for $r=2,4,\ldots,M$ when $M$ is even, and $r=2,4,\ldots,M-1$ when $M$ is odd. From these trigonometric moments, one can derive other characteristics of the NNTS axial distribution, such as the mean direction, circular variance, asymmetry coefficient, and kurtosis (see Mardia and Jupp, 2000).
In many practical applications, it is necessary to model axial data that are symmetric with respect to an angle $\mu\in(0,\pi]$, known as the axis of symmetry. A NNTS axial model is symmetric if the complex parameter vector $\underline{v}$ is restricted to be real (see Fernández-Durán and Gregorio-Domínguez, 2025). The corresponding density function is defined as:
$f_{NNTSaxialsym}(\theta\mid\underline{v}_{R})=\frac{1}{\pi}\left|\left|\sum_{k=0}^{M}v_{Rk}e^{i2k(\theta-\mu)}\right|\right|^{2}=\frac{1}{\pi}\sum_{k=0}^{M}\sum_{m=0}^{M}v_{Rk}v_{Rm}e^{i2(k-m)(\theta-\mu)}.$ (7)
Here, $\underline{v}_{R}$ is a vector of real coefficients. Alternatively, the complex vector $\underline{v}^{*}$ can be defined with $v_{k}^{*}=v_{Rk}e^{-i2k\mu}$. The total number of free parameters in a symmetric NNTS axial model is $M+1$.
@@
with $\underline{v}=(v_{0},v_{1},\ldots,v_{M})^{T}$ as the parameter vector and $\underline{e}=(1,e^{i2\theta},e^{i4\theta},\ldots,e^{i2M\theta})^{T}$ as the vector of trigonometric statistics, the estimation algorithm described by Fernández-Durán and Gregorio-Domínguez (2010) can be used. This involves a modified Newton algorithm on the complex unit hypersphere (manifold) $CS^{M}=\{\underline{v}:\sum_{k=0}^{M}||v_{k}||^{2}=1\}$. It is recommended to have at least $7M^{*}$ observations to reliably fit an NNTS axial model with $M=M^{*}$.
@@
where $\underline{\theta}$ is the vector of observed axial data, $\hat{\underline{v}}_{G}$ is the MLE under the general model, $\hat{\underline{v}}_{R}$ is the MLE under the restricted model, and $\hat{\ell}(\cdot)$ denotes the corresponding maximized log-likelihood values. Under standard regularity conditions, $LLR_{RG}$ asymptotically follows a chi-squared distribution with degrees of freedom equal to the number of constraints imposed to derive $\underline{v}_{R}$ from $\underline{v}_{G}$.
For instance, a test for symmetry can be performed by comparing the general NNTS axial model to a symmetric version. In this case, the test statistic follows a chi-squared distribution with $M-1$ degrees of freedom. To test for homogeneity across $P$ different populations with independent samples, one compares the sum of the maximized log-likelihoods from each population to the log-likelihood obtained from pooling all data into a single model. The degrees of freedom for the test are given by:
$\sum_{k=1}^{P}2M_{k}-2M,$
where $2M_{k}$ is the number of parameters in the fitted model for population $k$, and $2M$ is the number of parameters in the pooled model.
@@ 3.1 Feldspar Lath Orientations in Basalt
Feldspar crystals within rocks often align in parallel or subparallel patterns, providing valuable insights into the processes of rock formation and historical geological events, such as volcanic eruptions. Dataset B2 from Fisher (1993) contains measurements of the long-axis orientations of 133 feldspar laths in basalt. Table 1 (columns 2 to 4) presents maximized log-likelihood, Bayesian Information Criterion (BIC), and Akaike Information Criterion (AIC) values for fitted NNTS axial models with parameter $M$ ranging from 0 to 13. Figure 1 (left plot) shows a circular dot plot and histogram of the dataset along with the best-fitted NNTS axial densities selected by BIC ($M=0$) and AIC ($M=4$). A uniformity test by Fernández-Durán and Gregorio-Domínguez (2024), applied by doubling the axial angle, rejects the null hypothesis of uniformity implied by the BIC-selected model ($M=0$) in favor of the multimodal NNTS axial model chosen by AIC ($M=4$), yielding a p-value of less than 0.01.
Dataset B5 from Fisher (1993) contains orientations of an additional 60 feldspar laths in basalt. Fisher emphasized the importance of testing uniformity and determining the number of modes in this dataset. Table 1 (columns 5 to 7) presents maximized log-likelihood, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) values for NNTS axial models with parameter $M$ ranging from 0 to 6. Figure 1 (center plot) shows the histogram along with the best-fitting BIC and AIC models, both of which correspond to a unimodal and symmetric density with $M$=1. We conducted a likelihood ratio test comparing a symmetric NNTS model with $M=1$ against an asymmetric NNTS model with $M=2$, depicted as a dashed line in Figure 1 (center plot). The resulting p-value was 0.6077, indicating insufficient evidence to reject the simpler symmetric model ($M=1$) in favor of the asymmetric alternative ($M=2$). When comparing the symmetric and asymmetric NNTS axial models with ($M=2$) using a likelihood ratio test for symmetry, the null hypothesis of symmetry is rejected at a 5% significance level with a p-value equal to 0.0459.
Additionally, the uniformity test by Fernández-Durán and Gregorio-Domínguez (2024) produced a p-value between 0.01 and 0.05, leading to the rejection of the null hypothesis of uniformity at a 5% significance level when comparing the uniform density model ($M=0$) with the best AIC- and BIC-selected NNTS axial models with $M=1$. Our results are consistent with those of Fisher (1993), supporting both the non-uniformity and unimodality of the dataset.

### v2

@@ Abstract
The family of circular distributions based on non-negative trigonometric sums (NNTS), developed by Fernández-Durán (2004), is highly flexible for modeling datasets exhibiting multimodality and/or skewness. In this article, we extend the NNTS family to axial data by identifying conditions under which the original NNTS family is suitable for modeling undirected vectors. Since the estimation is performed using maximum likelihood, likelihood ratio tests are developed for characteristics of the density function such as uniformity and symmetry, as well as to compare different axial populations through homogeneity tests. The proposed methodology is applied to real datasets involving orientations of rocks, animals, and plants.
@@ 1 Introduction
Axial data, defined as data representing undirected vectors, are relevant in many applications across different disciplines, notably in Paleomagnetism for the statistical analysis of Earth’s magnetic field directions recorded during rock formation (Onstott, 1980; Tauxe, 2010). Axial data also appear in analyses of astronomical body orientations (Hawley & Peebles, 1975), protein structure axes (Barlow & Thornton, 1988), and certain anatomical structures in animals and humans (Vrtovec et al., 2009), such as the posterior corneal curvature of the eye (Koch et al., 2012). In Ecology, axial data assist in tracking animal magnetic orientation (Batschelet, 1981; Wiltschko & Wiltschko, 1995 and 2005; Fitak & Sönke, 2017) and in Environmental Science, they aid the study of vegetation orientation patterns, their relation to climate change (Sherratt, 2015), and associated effects on wind direction (Mingione et al., 2015).
Arnold & Sengupta (2011) describe various methods to generate densities for an axial random variable, $\theta\in[0,\pi)$, from either a linear random variable $X$ or a circular random variable $\phi\in[0,2\pi)$. The first method involves wrapping a linear random variable: $\theta=X\mod\pi$. The second, called adding, defines the axial density as the sum of a circular density at $\phi$ and $\phi+\pi$, equivalent to wrapping a circular variable: $\theta=\phi\mod\pi$. The third method uses polar transformation of a bivariate linear vector $(X,Y)$, defining $\theta=\tan^{-1}(Y/X)+\pi/2$. Early studies on the analysis of axial data initially approached the problem by transforming the axial variable $\theta$ into a circular random variable $\phi$ through a simple doubling transformation, $\phi=2\theta$ (Batschelet, 1981). The transformed data were then analyzed using standard models developed for circular data. This approach, commonly referred to as the method of angle doubling, produces models for axial data that are directly derived from corresponding models for circular data.
Common models for axial data include the axial von Mises (AvM), axial wrapped Cauchy (AWC), and sine-skewed distributions. The AvM model (Arnold & SenGupta, 2006), derived using the adding method, has the density:
@@
$f_{AWC}(\theta)=\frac{1-\rho^{4}}{\pi(1+\rho^{4}-2\rho^{2}\cos(2(\theta-\mu)))},\quad-\frac{\pi}{2}\leq\theta<\frac{\pi}{2},$
where $0\leq\rho\leq 1$ and $\mu$ is the location parameter. Additionally, Arnold & SenGupta (2011) introduced the angular central Gaussian (AACG) model, which generalizes the AWC by using a polar transformation of a bivariate Gaussian vector:
$f_{AACG}(\theta)\propto\left(1+\alpha\sin\left(2\left(\theta-\mu+\frac{\pi}{4}-\beta\right)\right)\right)^{-1},\quad 0\leq\theta<\pi,$
with parameters $\alpha$, $\beta$ and $\mu$. Abe et al. (2012) proposed sine-skewed (SineSk) axial models, introducing asymmetry through a perturbation term:
$f_{SineSk}(\theta)=(1+\lambda\sin(k(\theta-\mu)))f_{0}(\theta-\mu),\quad-\frac{\pi}{2}\leq\theta<\frac{\pi}{2},$
where $f_{0}(\theta)$ is the base symmetric axial density, $k=2$, and $1\leq\lambda\leq 1$. When $\lambda=0$, the SineSk model reduces to the symmetric base model. Multimodal axial densities may also be constructed as mixtures (e.g., mixtures of axial von Mises distributions), though this approach complicates numerical parameter estimation when the number of modes increases.
@@ 2.1 Definition
[…]
@@
[…]
where $c_{k}||^{2}=c_{re,k}^{2}+c_{im,k}^{2}$. To ensure identifiability of the $c$ parameters, we constrain $c_{0}$ to be a positive real number, which immediately implies that $c_{im,0}=0$. From Equation 2, it then follows that $c_{0}=\frac{1}{2\pi}-\sum_{k=0}^{M}\|c_{k}\|^{2}$.
@@
[…]
@@
[…]
This corresponds to doubling the original circular angle $\phi$, a common technique for analyzing axial data. For this to define a valid density, the parameters must satisfy
[…]
and, to ensure identifiability of the $v$ parameters, we constrain $v_{0}$ to be a positive real number. From Equation 5, it then follows that $v_{0}=\frac{1}{\pi}-\sum_{k=1}^{M}\|v_{k}\|^{2}$. Alternatively, the NNTS axial density can be expressed as:
@@
The number of free parameters in $\underline{v}$ equals $2M$, and $M$ determines the maximum number of modes for the NNTS axial density. The support of the axial random variable can be $(0,\pi]$ due to $f_{NNTSaxial}(\theta\mid\underline{v})=f_{NNTSaxial}(\theta\pm\pi\mid\underline{v})$, although the interval $\frac{\pi}{2},\frac{\pi}{2})$ is also common in literature. Note that NNTS axial models are nested; models with smaller $M$ are special cases of those with larger $M$, facilitating likelihood ratio tests. Specifically, the case $M=0$ corresponds to the uniform axial density.
The trigonometric moments, $E[e^{ir\theta}]$, for an odd $r$ are equal to zero. This is because, for a circular axial random variable,
$E[e^{ir\theta}]=E[e^{ir(\theta+\pi)}]=e^{ir\pi}E[e^{ir\theta}]=\cos(r\pi)E[e^{ir\theta}]=(-1)^{r}E[e^{ir\theta}],$
which equals $E[e^{ir\theta}]$ since $r$ is odd. Therefore, $E[e^{ir\theta}]=0$ for odd $r$. For even $r$, the trigonometric moments of the NNTS axial distribution in Equation (6) are given by
$E(e^{ir\theta})=\frac{1}{\pi}\int_{0}^{\pi}e^{ir\theta}\sum_{k=0}^{M}\sum_{m=0}^{M}v_{k}\bar{v}_{m}e^{i2(k-m)\theta}\,d\theta=\frac{1}{\pi}\sum_{k=0}^{M}\sum_{m=0}^{M}v_{k}\bar{v}_{m}\int_{0}^{\pi}e^{i(2(k-m)+r)\theta}\,d\theta.$ (7)
Since $k$ and $m$ are integers and $r$ is even, 2(km)+r2(k-m)+r is also even. Therefore, the integral $\int_{0}^{\pi}e^{i(2(k-m)+r)\theta}\,d\theta$ equals zero when $2(k-m)+r\neq 0$, and equals $\pi$ when 2(km)+r=02(k-m)+r=0. Thus, for even $r$, the trigonometric moment simplifies to:
$E(e^{ir\theta})=\sum_{m=\frac{r}{2}}^{M}v_{m-\frac{r}{2}}\bar{v}_{m}$ (8)
for $r=2,4,\ldots,M$ when $M$ is even, and $r=2,4,\ldots,M-1$ when $M$ is odd. From these trigonometric moments, one can derive other characteristics of the NNTS axial distribution, such as the mean direction, circular variance, asymmetry coefficient, and kurtosis (see Mardia & Jupp, 2000). Moreover, the invariance property of the maximum likelihood estimation method can be used to obtain estimates of functions of the parameters, such as the trigonometric moments and other derived quantities.
In many practical applications, it is necessary to model axial data that are symmetric with respect to an angle $\mu\in(0,\pi]$, known as the axis of symmetry. A NNTS axial model is symmetric if the complex parameter vector $\underline{v}$ is restricted to be real and nonnegative (see Fernández-Durán & Gregorio-Domínguez, 2025). The corresponding density function is defined as:
@@
Here, the vector $\mathbf{v}$ is restricted to be a nonnegative real vector, with the parameter space defined as $\{\mathbf{v}:v_{R0}>0,\;v_{Rk}\geq 0\text{ for }k=1,\ldots,M,\;\sum_{k=0}^{M}\|v_{Rk}\|^{2}=1\}$. Alternatively, the complex vector $\underline{v}^{*}$ can be defined with $v_{k}^{*}=v_{Rk}e^{-i2k\mu}$. The total number of free parameters in a symmetric NNTS axial model is $M+1$.
@@
with $\underline{v}=(v_{0},v_{1},\ldots,v_{M})^{T}$ denoting the parameter vector, $\underline{v}^{H}$ its Hermitian (conjugate transpose), $\underline{e}=(1,e^{i2\theta},e^{i4\theta},\ldots,e^{i2M\theta})^{T}$ the vector of trigonometric statistics, and $\underline{e}^{H}$ its Hermitian, the estimation algorithm described by Fernández-Durán & Gregorio-Domínguez (2010) can be applied. This algorithm involves a modified Newton method on the complex unit hypersphere parameter space (manifold), which can be locally approximated by a hyperplane known as the tangent space. The modified Newton algorithm of Fernández-Durán & Gregorio-Domínguez (2010) operates by searching for maxima in directions defined within the tangent space, using the complex gradient vector and the Hessian matrix of the log-likelihood function, and then reprojecting (retracting) onto the complex unit hypersphere.
For the symmetric NNTS axial model, at each iteration of the modified Newton algorithm, the norms of the elements of the complex vector $\mathbf{v}^{*}$ are computed, and the axis of symmetry, $\mu$, is optimized over a grid of values in the interval $0,\pi)$. After convergence of the modified Newton algorithm, the estimates of the nonnegative real parameter vector $\mathbf{v}$ and the axis of symmetry $\mu$ are obtained.
@@
[…]
For instance, a test for symmetry can be performed by comparing the general NNTS axial model to a symmetric version. In this case, the test statistic follows a chi-squared distribution with M1M-1 degrees of freedom. To test for homogeneity across $P$ different populations with independent samples, one compares the sum of the maximized log-likelihoods from each population to the log-likelihood obtained from pooling all data into a single model. The degrees of freedom for the test are given by:
$\sum_{k=1}^{P}2M_{k}-2M_{pooled},$
where $2M_{k}$ is the number of parameters in the fitted model for population $k$, and $2M_{pooled}$ is the number of parameters in the pooled model. Under the null hypothesis of homogeneity of the $P$ populations, and given the nested structure in the parameter $M$ of the NNTS axial models, we have $M_{k}=M_{\text{pooled}}$ for $k=1,2,\ldots,P$. Consequently, the number of degrees of freedom of the chi-squared likelihood ratio test for homogeneity simplifies to $2(P-1)M_{\text{pooled}}$. In practical applications of the homogeneity test, the researcher may know the value of $M_{\text{pooled}}$ from theoretical considerations or previous studies. If this is not the case, the researcher can apply the homogeneity test for different values of $M_{\text{pooled}}$ and calculate a corrected final $p$-value from the $p$-values of the tests corresponding to these different values. Various methods for correcting $p$-values due to multiple hypothesis testing can be used (Goeman & Solari, 2014; Cinar & Viechtbauer, 2022). In particular, we considered the Bonferroni and Tippett corrections for multiple testing, with the adjusted $p$-values defined as
$p_{c}^{(B)}=\min\left\{1,\min(p_{1},p_{2},\ldots,p_{P})\times P\right\}\quad\text{and}\quad p_{c}^{(T)}=1-(1-\min(p_{1},p_{2},\ldots,p_{P}))^{P},$
respectively, where $p_{c}^{(B)}$ and $p_{c}^{(T)}$ denote the Bonferroni- and Tippett-corrected $p$-values, and $p_{1},p_{2},\ldots,p_{P}$ are the individual $p$-values from the $P$ tests.
@@ 3.1 Feldspar Lath Orientations in Basalt
Feldspar crystals within rocks often align in parallel or subparallel patterns, providing valuable insights into the processes of rock formation and historical geological events, such as volcanic eruptions. Dataset B2 from Fisher (1993) contains measurements of the long-axis orientations of 133 feldspar laths in basalt. Table 1 (columns 2 to 4) presents maximized log-likelihood, Bayesian Information Criterion (BIC), and Akaike Information Criterion (AIC) values for fitted NNTS axial models with parameter $M$ ranging from 0 to 13. Figure 1 (left plot) shows a circular dot plot and histogram of the dataset along with the best-fitted NNTS axial densities selected by BIC ($M=0$) and AIC ($M=4$). A uniformity test by Fernández-Durán & Gregorio-Domínguez (2024), applied by doubling the axial angle, rejects the null hypothesis of uniformity implied by the BIC-selected model ($M=0$) in favor of the multimodal NNTS axial model chosen by AIC ($M=4$), yielding a $p$-value of less than 0.01.
Dataset B5 from Fisher (1993) contains orientations of an additional 60 feldspar laths in basalt. Fisher emphasized the importance of testing uniformity and determining the number of modes in this dataset. Table 1 (columns 5 to 7) presents maximized log-likelihood, Akaike Information Criterion (AIC), and Bayesian Information Criterion (BIC) values for NNTS axial models with parameter $M$ ranging from 0 to 6. Figure 1 (center plot) shows the histogram along with the best-fitting model according to BIC, which corresponds to a uniform axial density (M=M=0), and the best-fitting model according to AIC, which corresponds to a unimodal and symmetric density (M=M=1).
We conducted a likelihood ratio test comparing a symmetric NNTS model with $M=1$ against an asymmetric NNTS model with $M=2$, depicted as a dashed line in Figure 1 (center plot). The resulting $p$-value was 0.6077, indicating insufficient evidence to reject the simpler symmetric model ($M=1$) in favor of the asymmetric alternative ($M=2$).
Additionally, the uniformity test by Fernández-Durán & Gregorio-Domínguez (2024) produced a $p$-value between 0.01 and 0.05, leading to the rejection of the null hypothesis of uniformity at a 5% significance level when comparing the uniform density model ($M=0$) with the best AIC-selected NNTS axial model with $M=1$. Our results are consistent with those of Fisher (1993), supporting both the non-uniformity and unimodality of the dataset.
Table 2 presents the results of applying the NNTS axial homogeneity tests to the two populations of feldspar laths in basalt (with 133 from dataset B2 and 60 from dataset B5, respectively, both from Fisher, 1993) for values of $M_{\text{pooled}}$ ranging from 1 to 6. The table reports the maximized log-likelihood, BIC, and AIC values of the models fitted to the pooled population, as well as the $p$-values from the NNTS axial homogeneity tests. For $M_{\text{pooled}}=1,\ldots,6$, the null hypothesis of homogeneity is not rejected at the 5% significance level. The combined $p$-values, obtained by aggregating the six homogeneity tests over $M_{\text{pooled}}=1,\ldots,6$, are equal to 1.0000 under the Bonferroni correction and 0.6752 under the Tippett correction, indicating no evidence against the null hypothesis of homogeneity.